OpenStudy (anonymous):

If 3y+5=7x, then 21y-49x

OpenStudy (mathstudent55):

3y + 5 = 7x Subtract 7x from both sides, subtract 5 from both sides: 3y - 7x = -5 Multiply both sides by 7: 21y - 49x = -35